Two interfering sinusoidal waves are described by the wavefunctions: y1 = A sin(kx-ωt) and y2 = A sin(kx-ωt+φ). If φ cannot exceed π/2 (0 ≤ φ ≤ π/2), then which of the following is true about the resultant amplitude Ares?

1
Expert's answer
2021-12-14T15:58:27-0500

"y_1=A\sin(kx-\omega t),"

"y_2=A\sin(k x-\omega t+\varphi),"

"y=y_1+y_2=2A\cos\frac{\varphi}2\sin(k x-\omega t+\frac{\varphi}2),"

"A_{res}=2A\cos\frac{\varphi}2."